Question 5.)

Year Return
2016 13%
2017 13%
2018 -10%
2019 24%

part a.) What are the arithmetic mean and standard deviation of the historical returns above?

part b.) What is the geometric means of the returns above?

Please show formula and work
Answer #1

a.

Arithmetic Mean = (0.13 + 0.13 - 0.10 + 0.24)/4

Arithmetic Mean = 10.00%

Standard Deviationof Returns = 1/(4 – 1) (ri - 0.1021/2 i=1

Standard Deviation of Return = 0.1431

Standard Deviation of Return = 14.31%

b.

Geometric Return = [(1.13)(1.13)(1 - 0.10)(1.24)]1/4 - 1

Geometric Mean = 0.0926

Geometric Mean = 9.26%